a purely competitive firm finds that the market price for its product is $30 it has a fixed cost of $100 and a variable cost of $15 per unit for the first 50 units and then $35 per unit for all successive units. what is the average variable cost for the first 100 units? what output level will yield the largest possible profit for this purely competitive firm?
Answer #1

In order to calculate, Average Variable cost of first 100 units, We need the Total Variable cost of first 100 units.

The formula for calculating Average Variable cost is : Total Variable cost/ Quantity.

Given,

Variable cost for first 50 units = $ 15

And Variable cost for after every 50 units. = $35

Now, Total Variable cost of first 100 units

50 × $15 (Variable cost of first 50 units ) . + 50 × $ 35 (Variable cost after first 50 units)

= ( 50 × 15) + ( 50 × 35) = $2,500

Now, Average variable cost of first 100 units =

Total Variable cost of 100 units/ Quantity

= 2,500/100 = $25

b) The Profit maximising condition for Perfectly Competitive market is where Marginal Revenue / Price equals Marginal Cost.

In Perfect Competitive market, Price = Marginal Revenue.

If we consider the case of output ess than 50 units,

The Marginal Revenue/Price = $30 exceeds the Marginal cost = $ 15

And, In case of output more than 50 units,

The Marginal Revenue/Price= $30 is less than the Marginal Cost =. $35.

Thus, at the output level of 50 units ,the Profit will be maximised because in either left or right of 50 units, the Marginal Revenue ≈Marginal Cost